Solve for x (without using a calculator)
(log3)x2 = 2(log3)4 – 4(log3)5
All 3's are subscripts--base 3
The x2 means x squared

Answer by Alan3354(21583)

(log3)x2 = 2(log3)4 – 4(log3)5
log(x^2) = log(16) - log(625)
x^2 = 16/625
x = 4/25
No negative solutions are allowed due to log function.